OpenStudy (anonymous):

use a sum or difference identity to find exact value of sin 255.

OpenStudy (anonymous):

sin( 210 + 45 ) = sin 210 cos 45 + cos 210 sin 45

OpenStudy (anonymous):

sin 210 = sin ( 180 + 30 ) = sin180 cos 30 + cos 180 sin 30

OpenStudy (anonymous):

ok so where do i go from there?

OpenStudy (anonymous):

sin 210 = 0 . cos 30 + (-1) (1/2)

OpenStudy (anonymous):

sin 210 = - (1/2)

OpenStudy (anonymous):

then replace in first formula

OpenStudy (anonymous):

cos 210 = cos (180 + 30) = cos180cos30 - sin180sin30

OpenStudy (anonymous):

cos 210 = (-1) sqrt(3)/2 - 0. (1/2)

OpenStudy (anonymous):

cos 210 = - sqrt(3) / 2

OpenStudy (anonymous):

finally

OpenStudy (anonymous):